# Price formation without fuel costs: the interaction of elastic demand with storage bidding

### v1

@@ Abstract
Studies looking at electricity market designs for very high shares of wind and solar often conclude that the energy-only market will break down. Without fuel costs, it is said that there is nothing to set prices. Symptoms of breakdown include long phases of zero prices, scarcity prices too high to be politically acceptable, prices that collapse under small perturbations of capacities from the long-term equilibrium, cost recovery that is impossible due to low market values, high variability of revenue between different weather years, and difficulty operating long-term storage with limited foresight. We argue that all these problems are an artefact of modeling with perfectly inelastic demand. If short-term elasticity to reflect today’s flexible demand (-5%) is implemented in the model, these problems are significantly reduced. The combined interaction of demand willingness to pay and storage opportunity costs is enough to produce stable pricing. This behavior is illustrated by a model with wind, solar, batteries, and hydrogen-based storage, where the price duration curve is significantly smoothed with a piecewise linear demand curve. This removes high price peaks, reduces the fraction of zero-price hours from 90% to around 30%, and guarantees more price stability for perturbations of capacity and different weather years. Fuels derived from green hydrogen take over the role of fossil fuels as the backup of final resort. Furthermore, we show that with elastic demand, the long-term model exactly reproduces the prices of the short-term model with the same capacities. We then use insights from the long-term model to derive simple bidding strategies for storage so that we can also run the short-term model with limited operational foresight. We demonstrate this short-term operation in a model trained on 35 years of weather data and tested on another 35 years of unseen data. We conclude that the energy-only market can still play a key role in coordinating dispatch and investment in the future.
@@ 1.1 Problem statement
There is a widespread concern in the literature that if shares of wind and solar rise even higher, providing more than 80% of yearly electricity, prices might become so singular that the energy-only market might no longer function in a meaningful way. Taylor et al. [2] express the concern ‘because there are no fuel costs, employing nodal pricing as is would simply result in all prices being equal to zero’. Some simulations indicate that prices are not always zero, but most of the time, with very high prices in other hours. Mallapragada et al. [3] find that ‘in …VRE-dominant energy-only wholesale power markets …generators and storage facilities would earn the bulk of their annual energy market revenues in relatively few hours’. Similarly, Levin et al. [4] raise the concern that ‘…a large fraction of the value of [energy storage] may be realized during a relatively small number of periods’. The Royal Society [5] writes that ‘if paid only on the basis of short-run costs, the large-scale long-term storage …could never recover its capital costs’ and that ‘traditional spot markets, which were developed to deal with gas and coal powered generation, are not automatically …adaptable to …wind, solar, and storage.’
Additional concerns include an expectation of wide swings in prices between different years reflecting changing weather conditions [6], raising risks around cost recovery [7, 8, 9, 10], and concerns about how storage dispatch and prices behave when transitioning from long-term capacity expansion models with perfect foresight to short-term operational models with myopic foresight [11, 12]. The apparent challenge to secure sufficient revenue for assets needed for system reliability [13] has led to discussions about alternative market designs oriented towards capacity rather than energy provision [14, 15, 16].
@@ 1.2 Other literature
Other modeling studies incorporating newly-electrified sectors and demand flexibility in systems with high renewable shares have seen different behavior [17, 18, 19, 20]. These studies see more moderate price duration curves with fewer zero and scarcity prices in capacity expansion models that incorporated flexible demands from heating, transport, and coupling to hydrogen production, which frequently becomes price-setting. Similar effects have been observed for systems with storage [21]. Even without the dominance of renewables, it has been noted that a small amount of demand elasticity can improve market operation, screening away price spikes [22, 23, 24]. Others point out that systems without fuel costs might not undermine the working principles of energy-only markets if taking into account storage operation, scarcity prices and demand response [25].
@@
In power systems dominated by hydroelectricity, which also has near-zero marginal costs, it has been understood for decades how the opportunity costs for dispatch give the storage medium water a value [26, 27, 28, 29, 30, 31, 32, 33], thereby determining the bidding behavior of hydroelectric generators. These insights have been recently transferred to the dispatch of long-duration energy storage (like hydrogen storage or thermal energy storage in district heating), where marginal storage values have been used in operational dispatch settings to develop bidding rules [34] and deal with myopic foresight [35]. Bidding strategies for storage have also been investigated by Ward et al. [36, 37] without optimization methods, looking at how price variability affects arbitrage earnings of storage.
@@
A series of papers [38, 39] examines the price structure in a stylized setup with renewables and storage only and find that prices set by storage play an important role in the cost recovery of all assets. However, the analysis is simplified to a residual load duration curve setup without addressing intertemporal storage dynamics or storage capacity constraints. Furthermore, the authors assume inelastic demand and only one type of renewable generation and storage technology.
@@
While Antweiler and Müsgens [40] explore price setting and market equilibria using a study design with wind, solar, and storage technologies and demand response. The authors conclude that energy-only markets remain functional even when the system only includes generators with near-zero marginal costs. The analysis, however, does not consider the intertemporal dynamics of market operation or storage.
@@
Ekholm and Virasjoki [41] analyze pricing mechanisms in fully renewable systems with highly elastic demand (-25%) under perfect and Cournot-type imperfect competition and identify the key role of storage operation and demand elasticity for price setting. They consider a short-term operational setting with weekly time frames without capacity expansion, multi-decadal time series, or myopic foresight.
@@
De Jonghe et al. [42] describe different approaches to incorporate demand elasticity into capacity expansion models with perfect foresight, including cross-price elasticities. Among other ways, they provide a quadratic formulation based on inverse demand curves. In a model with high conventional capacity and no storage, they find that demand elasticity increases the share of wind and reduces peak capacity required as demand response clears the market in those peak hours. However, their analysis does not focus on the price formation mechanisms.
@@ 1.3 Main thesis
In this paper, we provide analytic insights and modeling results to provide a unifying framework for understanding price formation at high penetrations of variable renewables like wind and solar. We show how storage bidding and demand flexibility combine to stabilize price formation. We use a model complex enough to illustrate the salient points, yet simple enough to trace cause and effect.
@@
Our main contention is that much of the singular price behavior observed in the literature stems from the assumption of inelastic demand used in capacity expansion models for electricity. If more realistic demand behavior is used, with, for example, the -5% short-term price elasticity of demand at the average price that was recently observed from today’s industrial consumers in Germany [43, 44], then many artefacts resulting from assuming perfectly inelastic demand disappear. Prices no longer jump between zero and scarcity; they remain stable over many weather years and do not collapse with small capacity perturbations. This explains the smoother price results observed in sector-coupling studies that include demand flexibility (although not genuinely elastic demand) [17, 18, 20]. Furthermore, demand elasticity also smooths the marginal storage values for long-duration energy storage, which allows us to deduce heuristics for storage bidding based on these marginal values and dispatch the systems with myopic foresight similar to [35].
In addition to developing an understanding of these features theoretically through the interaction of shadow prices in the optimization model, we also provide hourly modeling simulations for three European countries featuring different renewable potentials (Germany, Spain, and the United Kingdom) with over 70 years of weather data. The present study is the first study we are aware of that considers elastic demand in such capacity expansion models with multi-decadal time series. Solving such models is enabled by innovations in interior-point solvers for such convex problems [45]. We illustrate price formation for various demand types and examine which demand modeling approaches cause the price formation mechanisms to diverge or align between long-term models with capacity expansion and short-term models with only dispatch variables. We then show how prices behave under capacity perturbations from the long-term equilibrium, how well simple storage bidding heuristics can reproduce the price formation with realistic myopic operational foresight, and how this affects cost recovery.
@@ 2.1 Introduction
Figure 1: Example market situations with demand curve ((1) and (2)) and different supply curves ((3) and (4)). Depending on the VRE availability illustrated for cases a) to d), either the demand (1), storage (2) and (4) or wind and solar production (3) sets the price.
We first provide some intuition by examining hourly supply and demand curves for a system with zero-marginal-cost variable renewable energy (VRE), a single infinite [^1] Scarcity and cost in energy storage capacity would introduce multiple price levels for charging and discharging bids over time (Fig. S1). storage technology and elastic demand. Fig. 1 shows how the price setting works for situations a) to d) with rising availability of VRE.
In the aggregated demand curve, part of the demand curve (1) is determined by the regular electricity demand, which is assumed to be elastic with a low magnitude (e.g. -5% at the average price). The other part is the willingness to pay for the storage charging (2). The storage willingness to pay is non-zero because the storage medium has a value. For hydrogen storage, this would correspond to the option to sell the hydrogen for a non-zero price.
@@
This scarcity pricing is important because it helps to set the opportunity costs of the storage, which can hold back energy for these high-priced hours. Because the demand curve is elastic, the storage constantly makes trade-offs for how long to hold its energy, store it, or discharge it. The opportunity costs set the MSV, which in turn determines the electricity prices in most hours through the storage bidding.
@@ 2.2 Mathematical theory
This section provides a mathematical representation of the price formation by mapping the prices to dual variables (or shadow prices) of an optimization problem. We first start with the long-term capacity expansion problem.
The long-run welfare maximization runs over electricity consumers $c$, generators $r$, and storage units $s$, each of which can have different costs and efficiencies. The optimization variables are the dispatch for each period $t$ of demand $d_{c,t}$, generation $g_{r,t}$, storage discharging $f_{s,t}$, charging $h_{s,t}$ and state of charge $e_{s,t}$ as well as capacities for each generator $G_{r}$, storage discharge $F_{s}$, charging $H_{s}$ and energy $E_{s}$.
@@
Specific annualized investment costs $c_{*}$ are applied for all assets, while linearised variable costs $o_{r}$ are given only for generators. We assume non-linear utility for the demand and linear bidding for generation, assuming constant short-term marginal generation costs.
@@
The price formation can be deduced from the stationarity equations of the Karush-Kuhn-Tucker (KKT) conditions. [^2] The KKT conditions are the first-order conditions necessary for an optimal solution. When the objective function to maximize is concave, and the constraints are affine, these conditions are also sufficient for optimality [46]. For the generation, we have
@@ 2.3 Conditions for matching prices in long-term and short-term models
In the long-term model (LT), storage and generation capacities are endogenous variables of the optimization problem. In this case, investment costs play a role in determining capacity investments that satisfy demand in a way that maximizes social welfare. In the short-term model (ST), where the capacities $G_{r},F_{s},H_{s},E_{s}$ are pre-determined and not part of the optimization. This raises the question of when prices in the LT and ST models can be the same, even though the objective function for the ST model only involves dispatch variables:
@@
By taking the optimal capacities $G_{r},F_{s},H_{s},E_{s}$ from the long-term model and setting them as constants in the short-term model, we can compare the resulting electricity prices $\lambda_{t}$.
We will now prove that the LT and ST model prices are identical as long as the aggregated inverse demand curve (aggregate of $U^{\prime}_{c,t}$) provides a unique mapping from the demand to the price through a strictly monotonically decreasing demand curve. If there are steps [^3] Steps would represent segments of perfectly elastic demand, e.g. load shedding for a fixed value of lost load (VOLL). in the shape of the demand curve, the prices in the LT and ST model can diverge.
@@
However, if the demand curve is a step function, such as with a VOLL at 2000 €/MWh, then prices are only identical during load shedding or excess supply hours in the absence of other generators with non-zero marginal cost. In the hours with load shedding, prices reach 2000 €/MWh in both the ST and LT models. In hours with excess supply beyond the storage charging capacity, prices are zero in both models. The constant demand for varying prices during non-shedding periods causes a demand-side degeneracy, which means the supply-side sets the price unless storage charging does. As investment costs can set the price in the LT model but not in the ST model, we no longer have identical prices.
If the supply curve were strictly monotonically increasing, e.g., with a variety of generators with different efficiencies and fuel costs, rather than in linear steps, this would also be sufficient to resolve the prices and guarantee identical prices between ST and LT models. Some structure is needed either on the supply or the demand to eliminate the degeneracy of prices.
3 Methods

### v2

# Price formation without fuel costs: the interaction of demand elasticity with storage bidding

@@ Abstract
Studies looking at electricity market designs for very high shares of wind and solar often conclude that the energy-only market will break down. Without fuel costs, it is said that there is nothing to set prices. Symptoms of breakdown include long phases of zero prices, scarcity prices too high to be politically acceptable, prices that collapse under small perturbations of capacities from the long-term equilibrium, cost recovery that is impossible due to low market values, high variability of revenue between different weather years, and difficulty operating long-term storage with limited foresight. We argue that all these problems are an artefact of modelling with perfectly inelastic demand. If short-term elasticity to reflect today’s flexible demand (-5%) is implemented in the model, these problems are significantly reduced. The combined interaction of demand willingness to pay and storage opportunity costs is enough to produce stable pricing. This behavior is illustrated by a model with wind, solar, batteries, and hydrogen-based storage, where the price duration curve is significantly smoothed with a piecewise linear demand curve. This removes high price peaks, reduces the fraction of zero-price hours from 90% to around 30%, and guarantees more price stability for perturbations of capacity and different weather years. Fuels derived from green hydrogen take over the role of fossil fuels as the backup of final resort. Furthermore, we show that with demand elasticity, the long-term optimisation model exactly reproduces the prices of the short-term model with the same capacities. We then use insights from the long-term model to derive simple bidding strategies for storage so that we can also run the short-term model with limited operational foresight. We demonstrate this short-term operation in a model optimised using 35 years of weather data and then tested on another 35 years of unseen data. We conclude that the energy-only market can still play a key role in coordinating dispatch and investment in the future.
@@ 1.1 Problem statement
There is a widespread concern in the literature that if shares of wind and solar rise even higher, providing more than 80% of yearly electricity, prices might become so singular that the energy-only market might no longer function in a meaningful way [2]. Taylor et al. [3] express the concern ‘because there are no fuel costs, employing nodal pricing as is would simply result in all prices being equal to zero’. Some simulations indicate that prices are not always zero, but zero most of the time, with very high prices in other hours. Mallapragada et al. [4] find that in ‘VRE-dominant energy-only wholesale power markets …generators and storage facilities would earn the bulk of their annual energy market revenues in relatively few hours’. Similarly, Levin et al. [5] raise the concern that ‘…a large fraction of the value of [energy storage] may be realised during a relatively small number of periods’. A study for the Royal Society [6] suggests that ‘if paid only on the basis of short-run costs, the large-scale long-term storage …could never recover its capital costs’ and that ‘traditional spot markets, which were developed to deal with gas and coal powered generation, are not automatically …adaptable to …wind, solar, and storage.’
Additional concerns include an expectation of wide swings in prices between different years reflecting changing weather conditions [7, 8, 9], raising risks around cost recovery [10, 11, 12, 13], and concerns about how storage dispatch and prices behave when transitioning from long-term capacity expansion models with perfect foresight to short-term operational models with myopic foresight [14, 15]. The apparent challenge to secure sufficient revenue for assets needed for system reliability [16] has led to discussions about alternative market designs oriented towards capacity rather than energy provision [17, 18, 19, 2].
@@ 1.2 Other literature
Other modelling studies incorporating newly-electrified sectors and demand flexibility in systems with high renewable shares have seen different behavior [20, 21, 22, 23]. These studies see more moderate price duration curves with fewer zero and scarcity prices in capacity expansion models that incorporated flexible demands from heating, transport, and coupling to hydrogen production, which frequently becomes price-setting. Similar effects have been observed for systems with storage [24]. Even without the dominance of renewables, it has been noted that a small amount of demand elasticity can improve market operation, screening away price spikes [25, 26, 27]. Others point out that systems without fuel costs might not undermine the working principles of energy-only markets if taking into account storage operation, scarcity prices and demand response [28].
@@
In power systems dominated by hydroelectricity, which also has near-zero marginal costs, it has been understood for decades how the opportunity costs for dispatch give the storage medium water a value [29, 30, 31, 32, 33, 34, 35, 36], thereby determining the bidding behavior of hydroelectric generators. These insights have been recently transferred to the dispatch of long-duration energy storage (like hydrogen storage or thermal energy storage in district heating), where marginal storage values have been used in operational dispatch settings to develop bidding rules [37] and deal with myopic foresight [38]. Bidding strategies for storage have also been investigated by Ward et al. [39, 40] without optimisation methods, looking at how price variability affects the arbitrage earnings of storage.
@@
A series of papers [41, 42] examines the price structure in a stylised setup with renewables and storage only and find that prices set by storage play an important role in the cost recovery of all assets. However, the analysis is simplified to a residual load duration curve setup without addressing intertemporal storage dynamics or storage capacity constraints. Furthermore, the authors assume perfectly inelastic demand and only one type of renewable generation and storage technology.
@@
Antweiler and Müsgens [43] explore price formation and market equilibria using analytical and empirical models with wind, solar, and storage technologies and demand response. The authors show that energy-only markets remain functional even when the system only includes generators with near-zero marginal costs, as long as free entry and competition ensure effective price setting. However, the analysis is limited to a stochastic state space covering only one hour, so that the intertemporal dynamics of storage operation is not modeled explicitly.
@@
Ekholm and Virasjoki [44] analyse pricing mechanisms in fully renewable systems with demand elasticity (-25%) under perfect and Cournot-type imperfect competition and identify the key role of storage operation and demand elasticity for price setting. They consider a short-term operational setting with weekly time frames without capacity expansion, multi-decadal time series, or myopic foresight.
@@
De Jonghe et al. [45] describe different approaches to incorporate demand elasticity into capacity expansion models with perfect foresight, including cross-price elasticities between different hours. In a model with high conventional capacity and no storage, they find that demand elasticity increases the share of wind and reduces peak capacity required as demand response clears the market in those peak hours. However, their analysis does not focus on the price formation mechanisms.
Adachi et al. [46] include price-responsive demand in a model with renewables and storage, and find that ‘stable average prices can be maintained throughout the transition to 100% VRE’. However, they did not consider multiple weather years, long-duration energy storage, or analyse the resulting price duration curves.
@@ 1.3 Main thesis
In this paper, we provide analytic insights and modelling results to provide a unifying framework for understanding price formation at high penetrations of variable renewables like wind and solar. We show how storage bidding and demand flexibility combine to stabilise price formation. We use a model complex enough to illustrate the salient points, yet simple enough to trace cause and effect.
@@
Our main contention is that much of the singular price behavior observed in the literature stems from the assumption of perfectly inelastic demand used in capacity expansion models for electricity. If more realistic demand behavior is used, with, for example, the -5% short-term price elasticity of demand at the average price that was recently observed from today’s industrial consumers in Germany [47, 48], then many artefacts resulting from assuming perfectly inelastic demand disappear. Prices no longer jump between zero and scarcity; they remain stable over many weather years and do not collapse with small capacity perturbations. This explains the smoother price results observed in sector-coupling studies that include demand flexibility (although this is often modelled with virtual storage units rather than explicit demand curves) [20, 21, 23]. Furthermore, demand elasticity also smooths the marginal storage values for long-duration energy storage, which allows us to deduce heuristics for storage bidding based on these marginal values and dispatch the systems with myopic foresight similar to [38].
In addition to developing an understanding of these features theoretically through the interaction of shadow prices in the optimisation model, we also provide hourly modelling simulations for three European countries featuring different renewable potentials (Germany, Spain, and the United Kingdom) with over 70 years of weather data. The present study is the first study we are aware of that considers demand elasticity in such capacity expansion models with multi-decadal time series. Solving such models is enabled by innovations in interior-point solvers for such convex problems [49]. We illustrate price formation for various demand types and examine which demand modelling approaches cause the price formation mechanisms to diverge or align between long-term models with capacity expansion and short-term models with only dispatch variables. We then show how prices behave under capacity perturbations from the long-term equilibrium, how well simple storage bidding heuristics can reproduce the price formation with realistic myopic operational foresight, and how this affects cost recovery.
@@ 2.1 Introduction
Figure 1: Example market situations with demand curve ((1) and (2)) and supply curves ((3) and (4)) for four different VRE feed-in situations. Depending on the VRE availability, either the demand (1), storage (2) and (4) or wind and solar production (3) sets the price.
We first provide some intuition by examining hourly supply and demand curves for a system with zero-marginal-cost variable renewable energy (VRE), a single infinite [^1] Scarcity and cost in energy storage capacity would introduce multiple price levels for charging and discharging bids over time (Fig. S2). storage technology and demand elasticity. Fig. 1 shows how the price setting works with rising availability of VRE.
In the aggregated demand curve, part of the demand curve (1) is determined by the regular electricity demand, which is assumed to have elasticity with a low magnitude (e.g. below -5% at the average price). The other part is the willingness to pay for the storage charging (2). The storage willingness to pay is non-zero because the storage medium has a value. For hydrogen storage, this would correspond to the option to sell the hydrogen for a non-zero price.
@@
This scarcity pricing is important because it helps to set the opportunity costs of the storage, which can hold back energy for these high-priced hours. Because the demand curve is sloped, the storage constantly makes trade-offs for how long to hold its energy, store it, or discharge it. The opportunity costs set the MSV, which in turn determines the electricity prices in most hours through the storage bidding.
@@ 2.2 Mathematical theory
This section provides a mathematical representation of the price formation by mapping the prices to dual variables (or shadow prices) of an optimisation problem. We first start with the long-term (LT) capacity expansion problem, in which the capacities of all assets are co-optimised with their dispatch. In the next section we look at the consequences of moving to a short-term (ST) model, where we fix all the capacities and optimise only the dispatch.
The long-run welfare maximisation runs over electricity consumers $c$, generators $r$, and storage units $s$, each of which can have different costs and efficiencies. The yearly welfare is represented by hourly weather and load conditions $t$. The optimisation variables are the dispatch for each period $t$ of demand $d_{c,t}$, generation $g_{r,t}$, storage discharging $f_{s,t}$, charging $h_{s,t}$ and state of charge $e_{s,t}$ as well as capacities for each generator $G_{r}$, storage discharge $F_{s}$, charging $H_{s}$ and energy $E_{s}$.
@@
Specific annual fixed costs $c_{*}$, made up of annualised investment costs and fixed operation and maintenance costs, are applied for all assets, while linearised variable costs $o_{r}$ are given only for generators. We assume non-linear utility for the demand and linear bidding for generation, assuming constant short-term marginal generation costs.
@@
The price formation can be deduced from the stationarity equations of the Karush-Kuhn-Tucker (KKT) conditions. [^2] The KKT conditions are the first-order conditions necessary for an optimal solution. When the objective function to maximise is concave, and the constraints are affine, these conditions are also sufficient for optimality [50]. For the generation, we have
@@ 2.3 Conditions for matching prices in long-term and short-term models
In the long-term model (LT), storage and generation capacities are endogenous variables of the optimisation problem. In this case, fixed costs play a role in determining capacity investments that satisfy demand in a way that maximises social welfare. In the short-term model (ST), the capacities $G_{r},F_{s},H_{s},E_{s}$ are fixed from the beginning and not part of the optimisation.
It is not obvious that there is a relationship between the hourly prices in the LT and ST optimisation models. Economic theory suggests they should converge, since if the ST model deviates from the LT equilibrium, asset owners should notice that they are making either losses or excess profits and then (in the long-term) exit or enter the market respectively. However, from the perspective of the optimisation models, they have different sets of variables and objective functions. The ST model doesn’t see the fixed costs of the assets at all, and the ST objective function only involves dispatch variables:
@@
By taking the optimal capacities $G_{r},F_{s},H_{s},E_{s}$ from the long-term model and fixing them as constants in the short-term model, we can compare the resulting electricity prices $\lambda_{t}$. Economic theory says the ST and LT should have the same prices, but we will show that in the optimisation models, this only occurs under certain conditions. As a result, prices from LT optimisation models should be treated with care.
The key point is that in the presence of storage the fixed costs from the LT objective, i.e. the terms multiplied with $c_{*}$ in (1), can mix themselves into the prices. This causes price deviations versus the ST model, where these fixed cost terms are not present. Without storage, the fixed costs are usually screened away in the screening curve analysis by a combination of the variable costs of the generators $o_{*}$ and the cost for load shedding in the form of the value of lost load (VOLL) [52]. With storage, the fixed costs for the storage play a critical role in setting the marginal storage values (MSV) $\lambda^{s}_{t}$, since the MSV enable the storage to recover their fixed costs from price arbitrage. The MSV then leach into the market prices whenever the storage is price setting. [^3] On a mathematical level, the fixed costs interact with the prices via the $\bar{\mu}$ variables in equations (3)-(6), which are then linked to the fixed costs via the stationarity equations for the capacity variables in the LT optimisation problem, e.g. for storage energy capacity $0=\frac{\partial\mathcal{L}}{\partial E_{s}}=-c^{e}_{s}+\sum_{t}\bar{\mu}^{e}_{s,t}$ [51].
We will now demonstrate that, in the presence of storage, the LT and ST model prices are identical as long as the aggregated inverse demand curve (aggregate of $U^{\prime}_{c,t}$) provides a unique mapping from the demand to the price through a strictly monotonically decreasing demand curve. If there are steps [^4] Steps would represent segments of perfectly elastic demand, e.g. load shedding for a fixed value of lost load (VOLL). in the shape of the demand curve, the prices in the LT and ST model can diverge.
@@
However, if the demand curve is a step function, such as with a VOLL at 2000 €/MWh, then prices are only identical during load shedding or excess supply hours in the absence of other generators with non-zero marginal cost. In the hours with load shedding, prices reach 2000 €/MWh in both the ST and LT models. In hours with excess supply beyond the storage charging capacity, prices are zero in both models. The constant demand for varying prices during non-shedding periods causes a demand-side degeneracy, which means the supply-side sets the price unless storage charging does. As fixed costs can set the price in the LT model but not in the ST model, we no longer have identical prices.
Symmetrically, if the supply curve were strictly monotonically increasing, e.g., with a variety of generators with different efficiencies and fuel costs, rather than in linear steps, this would also be sufficient to resolve the prices and guarantee identical prices between ST and LT models. Some structure is needed either on the supply or the demand to eliminate the degeneracy of prices.
3 Methods for simulations